A wave is modeled with the function y(x,t) = (0.46m) cos((0.72rad/m)x − (0.77rad/s)t + π/3). Find the period.

Round your answer to 2 decimal places


1
Expert's answer
2021-04-28T13:04:28-0400
"T=\frac{2\pi}{\omega}\\\lambda=\frac{2\pi}{0.77}=8.2\ s"